PULSE BURST LASER SYSTEM FOR HIGH SPEED LASER AND COMBUSTION MEASUREMENTS

\underline{NAIBO JIANG}, MATT WEBSTER, KATHRYN GABET, RANDY PATTON, JEFFEREY SUT-TON AND WALTER LEMPERT, Department of Mechanical Engineering, The Ohio State University, Columbus, OH 43210; JOSEPH MILLER AND TERRENCE MEYER, Department of Mechanical Engineering, Iowa State University, Ames, IA 50011; JENIFER INMAN, BRETT BATHEL, STEVE JONES AND PAUL DANEHY, NASA Langley Research Center, Hampton VA, 23681.

The critical features of a burst mode diagnostic imaging system is described, along with representative NO PLIF measurements at 1MHz, in a Mach 10 hypersonic flow, and Rayleigh imaging and CH PLIF at 10kHz in turbulent flames.